# Euclid Quick Data Release (Q1)

### v1

@@ 2.3 Noise scheduling
All models in this paper make use of the cosine-beta scheduler. The impact of this scheduler on the high dynamic range nature of astronomy data is explored in detail in Sect. 5.
2.4 Conditioning the model
Using the pipeline of DDPMs allows us to impose conditioning on our input space. Throughout training, the model is already conditioned on . Due to the cosine scheduler discussed previously, the model will alter an image very differently at a timestep close to , where the image is still very noisy, than when the image is close to timestep and near completion. However, we are not limited to only conditioning by timestep, often multiple parameters are provided alongside each image. Typically, this is in respect to ’class-conditioning’, where users can tune the expected output of the diffusion model to a specific class or category (Radford et al. 2021; Zhang et al. 2023). By providing an embedding space during training, the model learns not only the dynamics of the noise but the inherent differences of noise and non-noise pixel values for each of the representations. This conditioning provides small nudges to the random walk, guiding the denoising to the respective population of the class. Providing rich representations to the model has led to the impressive text-to-image capabilities of popular diffusion-based implementations, such as Dall-E (Ramesh et al. 2022) or Imagen (Saharia et al. 2022).
This type of conditioning allows the user to direct the image generation to a particular class or style of image. However, with the possibility of complex prompts and classes, it may be difficult to know exactly what the output image will look like. Furthermore, with diffusion’s ability to cover large areas of the search space, a repeat sample of the same prompt is likely to produce an image very different from the first, even if the generated image belongs to the requested class. This is due to the one-to-many relationship between classes/prompts and their subsequent generated outputs.
In a situation where the user requires consistency in part of the image, a more direct manipulation of the produced pixels is required. A process known as ’inpainting’ (Pathak et al. 2016; Yeh et al. 2017; Yu et al. 2018; Wang et al. 2018), forces conditioning of specific pixels within the image, allowing the model to fix certain parts of the image whilst generating others – all whilst remaining coherent and consistent across the image as a whole. Allowing the user to selectively remove and replace certain aspects of images has gained a lot of popularity in the last few years due to its mainstream adoption by nearly all smartphone brands (Sargsyan et al. 2023).
2.5 Repainting
This paper uses a specific implementation of inpainting using diffusion known as Repaint (Lugmayr et al. 2022). Due to the model being conditioned on timestep , where it is expecting the whole image to be at a certain noise level, we can not simply take an image and a corresponding mask and replace the masked pixels with noise. The model itself does not know that a mask has been applied nor does it know that the inverse mask should remain unaffected. Instead, it will continue to remove noise across the whole image as usual, leading to the clean parts of the image becoming corrupted. It is also not possible to generate an entirely new image and copy over the masked pixels to the original image as there will be no continuity between the existing and new elements.
To combine the parts of the image we want unchanged, and the parts we want to generate, both the forward pass and the backward pass of the diffusion process must be utilised. This pipeline needs to generate an image that incorporates the mask and the fixed pixels for each timestep , as shown in Fig. 1.
Input
The main input of each timestep of the repaint pipeline is the denoised image, , from the previous timestep. To aid in the understanding of the pipeline, this image is referred to as for this discussion. At the first iteration when , the input is simply an image of pure Gaussian noise.
Fixed pixels
To ensure that the pixels we want unchanged are still present at the end of the diffusion process, we must interject these fixed pixels during each timestep. However, as the model is expecting all pixels to be at a specific noise level, as described in Eq. 2, we must utilise the same scheduler used in training to add the correct amount of noise for the current timestep . Due to the stochasticity present in adding Gaussian noise to the image, the pipeline allows us to create different versions of the input with the required noise. By averaging these samples, the important details of the image are more consistently preserved, ensuring that once all iterations are complete, the fixed pixels look identical to the original input. We refer to the image formed from the average of the N samples as .
Utilising the mask
To combine the parts of the image we want to keep and the parts we want to generate, we utilise a mask , which is an array of equal size to our image, with pixel values of either zero or one. Pixels valued as one will be generated from scratch by the diffusion process, with pixels valued zero being replaced at each noise step with the fixed pixels in from the process described above. The two parts can then be combined into a coherent image and then passed back into the diffusion model to be further denoised and progress onto timestep ,
By denoising the combination of pixels we want to keep with the generated pixels, we condition the diffusion process to ensure that in the final image, all pixels are coherent, being conditioned on the details found within the fixed pixels.
Incorporating the Repaint method for AGN identification has the following benefits
Diffusion methods have been proven their ability in many applications to accurately recreate complex data sets, allowing for realistic generation of galaxy morphology images.
There are no constraints on the size and shape of masks used and, since Repainting is only applied at inference, no additional retraining is required.
There exists a large imbalance of normal galaxies compared to AGN, allowing for the model to learn a bias. This can be exploited to create an outlier detection-based classifier without explicitly requiring labels during training.
@@ 3.1 Sample selection
Covering 63.1 deg2, the Euclid Q1 Data Release (Euclid Quick Release Q1 2025) contains just under 30 million catalogued sources (Euclid Collaboration: Romelli et al. 2025) in the Euclid Deep Fields North (EDF-N), South (EDF-S), and Fornax (EDF-F). In this work we use exclusively the VIS images (Euclid band), because we aim to identify AGN candidates on a single image. However, due to the varying distances and sizes of sources, an appropriate subset must be chosen to ensure the model is presented with a sufficiently broad selection of resolved galaxies. The following criteria are applied to the 30 million source catalogue of Q1:
@@
the source must have low probability of being spurious (spurious_prob_euclid 0.2);
@@
After this sample selection we are left with 1 142 606 sources across the three Q1 fields. VIS image cutouts (6464 pixels) were created for each of the respective sources, where problematic images, such as those at the very edge of tiles where a full cutout could not be created, were removed. This provided us with images and catalogue data for 1 060 864 sources. We then apply a 80 : 20 train and test split. Even though the inpainting and its results are a separate pipeline to the diffusion training, to prevent any potential data leakage, all of the decision boundaries created are based solely on the training set. When showing classifier performance, all results are from the unseen 20% of the data. This is to prevent any potential memorisation of the images from training impacting the results. However, in our analysis, no obvious differences have been found in the inpainting results between the two sets, showing that the model is not simply memorising the training set.
@@
the source must have a low probability of being star. (phz_star_prob_phz_class 0.1); and
@@ 3.2 Comparisons with other classifications
Although the training of the diffusion model does not make use of any labels, it is important to compare the distributions of metrics over different source selection criteria. This allows us to compare both the performance of detecting particular sources but also to determine what sources or factors may hinder the pipeline’s ability to classify objects sufficiently.
In particular, we use some of the sample definitions discussed in Euclid Collaboration: Matamoro Zatarain et al. (2025). A selection of stars are used from DESI (DESI Collaboration et al. 2024) and Gaia (Gaia Collaboration: Vallenari et al. 2023; Gaia Collaboration: Bailer-Jones et al. 2023). The normal galaxies sample is also from DESI (DESI Collaboration et al. 2024). Comparisons are made to selections of AGN and quasars using Euclid photometry with the _ and _ selections from Euclid Collaboration: Matamoro Zatarain et al. (2025) and the 2-colour selection from Euclid Collaboration: Bisigello et al. (2024, hereafter Bisigello24-A) . Additional AGN and QSO selections are used from DESI (DESI Collaboration et al. 2024), Gaia/Quaia (Storey-Fisher et al. 2024), as well as the 90% reliability WISE sample from (Assef et al. 2018), which is referred to as R90 for the remainder of the paper.
As is particularly true for AGN and QSO, it is impossible to create a selection criterion that will identify both with high completeness and high purity a single population of objects. Instead, we will use the aforementioned samples to assess contamination and purity of our method to the extent that is possible.
@@ 3.3 Data expectations
Although, by definition, a difference between a galaxy with and without a bright central component should be found, it is important to validate that this behaviour can be measured with the Euclid VIS images. To quantify this difference we analyse the relationship between the brightest central pixel and its surrounding pixels. By producing ratios comparing the central pixel to 1 and 2 pixels wide regions around the central pixel, we form a rate of change (ROC) metric over sources, as described by
@@
Pixels surrounding the brightest pixel, which share similar values to the maximum will produce a ROC value closer to zero, implying that the source does not feature a prominent active component. Sources featuring a single very bright pixel, surrounded by fainter pixels will have a ROC value closer to its maximum of 2. The produced distributions of sources, analysed according to both magnitude and various selection criteria (Sect. 3.2), are shown in Fig. 3. It is evident that sources with a high flux point-spread function (PSF) light profile show the highest ROC values, as expected. This effect is seen in the star sample and the QSO dominated selections (R90 WISE, Quaia, DESI-QSO, and Euclid-QSO). On the other hand, spectroscopically confirmed normal galaxies show very broad ROC distributions.
@@ 3.4 Addressing the use of asinh for image scaling
Given its prevalence in galaxy morphology-based research (Lupton et al. 2004), the effect of using an transformation on the data is also investigated. The function, which is often to the benefit of visual inspections, reduces the differences between the brightest and faintest pixels. However, when attempting to differentiate galaxy and non-galaxy sources, the function becomes a hindrance, causing the distributions to more significantly overlap, as shown in Fig. 4. To generate a measure for class overlap we measure the ROC for each image, taking the mean of each non-galaxy class. We use the percentage of our true galaxy sample that have a ROC value over the respective class mean as our overlap metric. The extent of the overlap is showcased in Table 1, where galaxy overlap increases by at least a factor of 2.5 and in the extreme case, nearly 50 times more galaxies overlap with stars. In the following sections of the paper we will discuss the set up, training and performance of the diffusion model that automatically learn to reconstruct the central emission expected for the normal galaxy population, and thus by contrast allow us to identify stars and AGN.
4 Training and inference pipeline
The diffusion and inpainting methods are two distinct pipelines, one for the initial training of the model and a second for the actual inpainting. Training the diffusion model makes use of the whole image and learns to predict what noise to remove at each timestep. This allows for the generation of novel examples from the trained distribution.
4.1 Training the diffusion model
For training, we use 6464 pixel VIS cutouts for each of the selected sources discussed in Sect. 3.1. We use the training pipeline described in Dhariwal & Nichol (2021)111Available at https://github.com/openai/guided-diffusion with the loss function
The hybrid loss is a combination of the mean squared error (MSE) and the variational lower bound (VLB). Using MSE ensures the output pixels of the model are numerically similar to the original image, maintaining structural and visual consistency within the generated images.
The first term in 6 is
where and are the true and predicted noise added for a timestep, respectively.
The VLB loss, shown in Eqs. 8 to 10, measures how well the diffusion process has modelled the underlying distribution of the image data, ensuring that the image as a whole is coherent and plausible for the data. The VLB loss has a weighting applied so as to not overpower the MSE loss:
By utilising the Kullback-Leibler (KL) divergence (Kullback & Leibler 1951), a measure of the difference of two distributions, the loss acts as a regulariser to prevent simple memorisation and instead optimising for learning general features and patterns present in real data.
By combining both losses, the model becomes accurate at generating realistic data at both the micro (pixel-wise) and macro (image-wise) levels. This is vital in ensuring the model learns the dynamics of morphology images through the diffusion process, allowing it to effectively recreate essential features through repainting. Throughout the rest of the paper the loss presented in Eq. 6 is referred to as the default loss.
Alternative loss
To address this, we applied a normalisation to by dividing each image’s error by its respective maximum pixel value, which could be from a brighter external source. This allows for the errors of both bright and faint images to scaled more equitably. The output of this loss is still combined with errors calculated in . Throughout the remainder of the paper, we refer to this variant of the loss as the normalised loss.
To visualise the impact of amending the loss function, with respect to an images brightness, Fig. 5 shows how the error from the unmodified MSE changes as the model is trained for longer. Each line represents the error for a collection of images in a random batch, with the increased transparency indicating the passage of time where the model had been trained on more epochs. The final epoch of each model is presented by a black line. The effect of removing point-like sources, as discussed in Sect. 3.1, is also explored.
An initial observation that can be made with this representation is the distribution of the number of images of a particular brightness. By focusing solely on the black lines, where the impact is most noticeable, each vertex on the line represents an individual image. The more frequent updates the line has within a range of maximum pixel values, the more images in our data set exist within that region. As the line has more updates when the max pixel value is less than one, there are significantly more images with fainter pixels. This further highlights the disproportionate penalisation of MSE as it assigns the majority of the error to a minority of the data.
The two models which feature the reduced training set (bottom row) share similar dynamics during training, regardless of the loss used. This is largely due to many point-like sources also being some of the brightest images and therefore the scale of the data set is reduced. As the most problematic data are removed, the impact of disproportionate penalisation is significantly reduced, allowing the fainter sources to be the focus during training. After analysing the results from each of the four models, which will be discussed in Sect. 6.2, we chose to use the model trained on the full data set using the normalised loss.
4.2 Mask creation
Due to the nature of the application, our aim is to mask out any suspected AGN without the need to rely on labels ahead of time. Each mask is positioned so that brightest pixel in the centre of the source becomes the centre of the mask. Although each image is centred on the source, various interactions and dynamics allow for suspected AGN to be offset from the centre. To allow for some deviation, a 99 pixel window around the centre of the image is used to detect the brightest pixel. This offset allows for the correct masking of less symmetric sources whilst limiting the likelihood of masking an adjacent source by mistake. The masks for each image are assigned dynamically throughout the data loading pipeline.
Throughout this paper we use a 55 pixel mask, unless explicitly stated otherwise, centred on the calculated brightest pixel. A 55 pixel mask, which covers 0505, allows us to effectively cover the AGN contribution whilst minimising loss of the surrounding galaxy structure. Given that the VIS PSF has a full width half maximum (FWHM) of 013 (Euclid Collaboration: Mellier et al. 2024), the AGN’s light is spread over multiple pixels. Taking into account the redshift of the galaxy, AGN within nearby sources () will cover 4.8 pixels, ensuring that we can fully encapsulate the core without unnecessary overlap of the host morphology. By retaining as much as the host galaxy information as possible for inpainting, we also reduce the risk of introducing artificial features into the reconstructed image.
4.3 Reconstruction rescaling
An important difference with the processing of data in this work compared to other morphology-based papers is how the data input into the network are the raw values direct from the sensor, rather than images binned to the standard 0–255 pixel scale range. As comparisons of individual pixels are being made on such a small mask, the relative differences become much more impactful and would likely be lost if values were binned.
One potential issue with this setup is that the libraries used for training are optimised for images within the standard pixel range. Although this did not prevent the model from effective learning the distribution, it does require a rescaling of the generated output as the range is restricted to between 1 and 1 to improve training stability.
We experimented with simply normalising the input data so that all our images were in the range 0–1; however, due to the large dynamic range of the data, the faintest pixels ended up being . Given the previous discussions on how MSE does not penalise smaller errors effectively, the model could not accurately recreate sources, outputting a near empty image of very low pixel values.
Instead, to allow the comparison of the original and outputted pixel values, we make use of the fact that 99.5% of the image (4071 of 4096 pixels when using a 55 pixel mask) are shared between the input and output. The following scaling can be dynamically applied using each images respective values to allow the non-masked pixels to be consistent:
where
and
5 Background noise versus diffusion noise
It is important for clarification to address the distinctions between the definition and use of ‘noise’ in diffusion models compared to the traditional understanding of noise in astronomy-based data. In the context of astronomical imaging, noise refers to the random fluctuations inherent in observational data, often caused by e.g. background emission, electronics read-out noise, and atmospheric disturbances. This ‘background noise’ reduces the signal-to-noise ratio (S/N) of the image, affecting the ability to detect subtle details in any analysis.
In contrast, the diffusion process uses an intentional integration of Gaussian noise to facilitate training of the generative model. The pixels of image , follow the distribution: , rather than following the distribution of the background noise. It is unclear whether using the background noise as the pure noise sample would cause issues for the model. Given the aim of the diffusion process is to denoise a noisy image so that a realistic one emerges, the original noise should be largely, if not entirely, removed. Given the background noise constitutes any area that is not significantly affected by a source, the model may attempt to remove it which would create an unrealistic image. On the other hand, it may result in less pixels needing to be denoised, optimising the models ability at learning and processing the source features. Such an experiment is left for future exploration.
5.1 The effect of signal-to-noise
In machine-learning research, the improvements of generative models are often demonstrated with data sets that are of high resolution and typically noise free. This provides a level of consistency across the full data set, ensuring the benchmarks reflect the model’s performance under ideal conditions. Whether through loss functions (as discussed in Sect. 4) or in the practical application of the pipelines, how the model behaves on one part of the data will be similar to another.
However, the difficulty in learning the dynamics of our data are not consistent throughout the images. Higher signal-to-noise ratios and lower magnitudes provide higher contrast between the source features and the background noise, enabling the model to discern these features more accurately and faster than the low S/N counterparts. In the context of diffusion models, where the aim throughout training is to learn the dynamics of noise at different timesteps, the varying S/N of the images could again cause inconsistencies during training.
Figure 6 explores how this variability in data impacts the scheduled noise added to the data. By measuring the specific pixels that are dominantly background noise (top) and those that are significantly affect by the source (bottom), and further splitting images into S/N bins, we show the transformation from raw pixels to Gaussian noise across the entire scale of our data.
Analysing the top row, we can see that low S/N images are more significantly affected by the introduction of noise. This also has an effect much earlier on in the scheduler. The background noise converges to the diffusion’s pure Gaussian distribution, across all S/N bins, at between 1/2 () and 3/4 () of the total noise steps. When focusing on the pixels dominated by the source (bottom row), high S/N images are fairly unaffected for large parts of the schedule, with the brightest images not changing significantly until half way (). It is not until the scheduler is within the final 250 steps that the brightest images converge to the Gaussian distribution.
To visualise the true effect of this differing behaviour across S/N, Fig. 7 shows a sample of images throughout different timesteps. Each image shown is a sample from the respective S/N bins below. It is clear that the noise scheduling does not perform uniformly across the images.
To increase performance of a diffusion model, whether to recreate a data set more accurately or increase the complexity of the data you are modelling, the simplest way is to increase the number of unique denoising steps performed. The important part here is the requirement to be unique, as simply adding more timesteps often results in the additional timesteps (especially in the latter stages) looking too similar and therefore become just additional overhead. This is the exact reason Nichol & Dhariwal (2021) implemented the cosine-beta scheduler, discussed Sect. 2.3, the original linear scheduler produced images that were pure noise too early into the process.
The images in Fig. 7 show that there are further optimisations that could be made to improve the scheduler for our data set. For optimal performance, one should expect the entire grid of images to behave like those in final 2 or 3 columns. A scheduler that adapts to the specific S/N of the image would be required and is left for future work. As our work is specifically using the reconstruction loss of inpainting as a classifier, the current scheduler is sufficient for our science goals.
@@ 6 Results
To explore the ability of the model at recreating galaxy images, we apply a variety of different masks to a selection of unseen images. Figure 8 shows the original image on the leftmost column, followed by the regenerated images in subsequent columns. With each mask, shown on the top row, the white area shows the pixels that are fixed from the original image whereas the black areas are generated using the diffusion model.
As our model’s aim is not to generate large collections of synthetic data, the variety and scientific validity of the produced galaxies are not put under high scrutiny, instead we only check that there are no obvious catastrophic mistakes. In the last two columns, we can see that due to the large mask and fewer pixels to guide the reconstruction, some of the images produce a visible boundary around the image where the combination of fixed pixels and generated pixels are not as seamless as the other images. Given the masks used for AGN detection are significantly smaller than those presented in Fig. 8, the fraction of pixels that the inpainting is conditioned on will be much higher, providing tighter constraints on the output compared to these examples.

### v2

@@ 3.2 Reconstruction rescaling
and
@@ 4.1 Sample selection
Covering 63.1 deg2, the Euclid Q1 Data Release (Euclid Quick Release Q1 2025) contains just under 30 million catalogued sources (Euclid Collaboration: Romelli et al. 2025) in the Euclid Deep Fields North (EDF-N), South (EDF-S), and Fornax (EDF-F). In this work, we use exclusively the VIS images (Euclid band) because we aim to identify AGN candidates on a single image. However, due to the varying distances and sizes of sources, an appropriate subset must be chosen to ensure the model is presented with a sufficiently broad selection of resolved galaxies. The following criteria are applied to the 30 million source catalogue of Q1:
@@
the source must have a low probability of being spurious (spurious_prob_euclid 0.2);
@@
It is crucial to note that because AGN are intrinsically rare compared to the general galaxy population, the training set does not need to be pre-filtered to remove them. The diffusion model, in learning the data distribution, naturally develops a prior that is heavily biased towards the ‘typical’ light profiles of non-active galaxies. This inherent bias is the key mechanism that enables the subsequent outlier detection.
After this sample selection, we are left with 1 142 606 sources across the three Q1 fields. VIS image cutouts (6464 pixels) were created for each source. Problematic images, such as those at the very edge of tiles where a full cutout could not be created, were removed. This provided us with images and catalogue data for 1 060 864 sources. We then apply a 80 : 20 train and test split. Even though the inpainting and its results are a separate pipeline to the diffusion training, all of the decision boundaries created are based solely on the training set to prevent any potential data leakage. When showing classifier performance, all results are from the unseen 20% of the data. This prevents any potential memorisation of the training images from impacting the results. However, in our analysis, no obvious differences have been found in the inpainting results between the two sets, showing that the model is not simply memorising the training set.
@@
The source must have a low probability of being a star. (phz_star_prob_phz_class 0.1); and
@@ 4.2 Comparisons with other classifications
Although our diffusion model is trained in an unsupervised manner without labels, it is essential to compare the distributions of metrics over different source selection criteria. This allows us to compare the performance of detecting particular sources and determine what sources or factors may hinder the pipeline’s ability to classify objects sufficiently.
In particular, we use some of the sample definitions discussed in Euclid Collaboration: Matamoro Zatarain et al. (2025). A selection of stars are used from DESI (DESI Collaboration et al. 2024) and Gaia (Gaia Collaboration: Vallenari et al. 2023; Gaia Collaboration: Bailer-Jones et al. 2023). The normal galaxies sample is also from DESI (DESI Collaboration et al. 2024). Comparisons are made to selections of AGN and quasars using Euclid photometry with the _ and _ selections from Euclid Collaboration: Matamoro Zatarain et al. (2025) and the 2-colour selection from Euclid Collaboration: Bisigello et al. (2024, hereafter Bisigello24-A) . Additional AGN and QSO selections are used from DESI (DESI Collaboration et al. 2024), Gaia/Quaia (Storey-Fisher et al. 2024), as well as the 90% reliability WISE sample from (Assef et al. 2018), which is referred to as R90 for the remainder of the paper. The reader is directed to Euclid Collaboration: Matamoro Zatarain et al. (2025) for a comprehensive overview of each of the selections used throughout our paper.
In this work, when referring to completeness against a specific reference sample (e.g., a catalog of selected AGN candidates), we mean the fraction of objects in that reference sample that are successfully identified by our method. This is due to the true astrophysical completeness being unknown and extremely difficult to measure from the observational data alone. Our context is equivalent to the standard machine learning metric of recall for the positive class, which is defined in Eq. 13.
When developing a method for initial candidate selection, where the goal is to flag objects for follow-up or multi-wavelength analysis, high recall is often prioritised. The aim is to minimise the loss of potentially interesting or rare sources (i.e., reduce false negatives), even if this comes at the cost of lower precision (i.e., including a larger fraction of contaminants or false positives in the selected sample). These contaminants can often be filtered out through subsequent analysis stages or by incorporating additional data. Our primary goal here is to demonstrate the method’s capability to recover a wide variety of known AGN and QSO types, making recall a key indicator of its potential utility in broad searches.
As is particularly true for AGN and QSO, it is impossible to create a selection criterion that will identify a single population of objects with high completeness and high purity. Instead, we will use the aforementioned samples, many of which are only candidates themselves, to assess the contamination and purity of our method to the extent that is possible.
@@ 4.3 Data expectations
As a baseline for our morphology-based approach, it is first necessary to verify that the Euclid VIS images contain sufficient information to resolve the contrast between the compact central emission characteristic of an AGN and the broader light profiles of typical galaxies. We analyse the relationship between the brightest central pixel and its surrounding pixels to quantify a difference between normal galaxies and AGN. By producing ratios comparing the central pixel to 1 and 2-pixels-wide regions around the central pixel, we form a rate of change (ROC) metric over sources, as described by
@@
Pixels surrounding the brightest pixel, which share values similar to the maximum, will produce a ROC value closer to zero, implying that the source does not feature a prominent active component. Sources featuring a single very bright pixel surrounded by fainter pixels will have a ROC value closer to its maximum of 2. The produced distributions of sources, analysed according to both magnitude and various selection criteria (Sect. 4.2), are shown in Fig. 3. As expected, sources with a high flux point-spread function (PSF) light profile show the highest ROC values. This effect is seen in the star sample and the QSO-dominated selections (R90 WISE, Quaia, DESI-QSO, and Euclid-QSO). On the other hand, spectroscopically confirmed normal galaxies show very broad ROC distributions.
@@ 4.4 Addressing the use of asinh for image scaling
Given its prevalence in galaxy morphology-based research (Lupton et al. 2004), the effect of using an transformation on the data is also investigated. The function is often used to benefit visual inspections by reducing the difference between the brightest and faintest pixels. In our approach, however, when attempting to differentiate galaxy and non-galaxy sources, the function becomes a hindrance, causing the distributions to more significantly overlap, as shown in Fig. 4. To generate a measure for class overlap, we measure the ROC for each image, taking the mean of each non-galaxy class. We use the percentage of our true galaxy sample with a ROC value over the respective class mean as our overlap metric. The extent of the overlap is showcased in Table 1, where galaxy overlap increases by at least a factor of 2.5 and in the extreme case, nearly 50 times more galaxies overlap with stars.
In the following sections of the paper, we will discuss the results of the diffusion model and analyse the performance of the inpainting pipeline that automatically learns to reconstruct the central emission expected for the normal galaxy population, allowing us to identify stars and AGN as outliers.
@@ 5 Results
Although the focus of the paper is on the model’s performance on small areas of a galaxy’s core, Repaint’s adaptability allows the pipeline to handle a variety of mask shapes and sizes. More general examples of the inpainting performance can be seen in Fig. 17. With larger masks, fewer pixels remain for conditioning, making the output more susceptible to artifacts. However, Lugmayr et al. (2022) demonstrate that this can be mitigated by increasing the number of resamples during the Repaint loop, at the cost of longer inference times.